OpenStudy (anonymous):

x+5/x=x+1/x+2=? A-5 B-2.5 C-7.5 D-10 HELLPPP!!

OpenStudy (anonymous):

I Can't See The Picture Because I'm Not A Student Of Connexus........

OpenStudy (anonymous):



OpenStudy (anonymous):

you give out 4 options to choose or can i have other solution?

OpenStudy (anonymous):

Let me try to solve this. In my opinion, you have to solve for x first and then put the value of x into the fraction to get the solution. that means: x +5 /x = X + 1 / x -2. since x>0, we can calculate the equation like ( (x-2) (x+5) = (x +1) x . --->x^2 +3x -10 = x^2 + x ---> 4x = 10 ---> x = 5/2. Now, we have x, replace to x+5 /x to know the scale. It is (5/2 +5)/ 5 /2= 3. conclusion: x +5 / x = 3 hope this right